OpenStudy (anonymous):

Can someone help me with this question: The amount of cocaine in a person's body fluids is modelled by the differential equation: y'(t)= -ky(t)/(a+y(t) where A,k are positive constants. For cocaine, A is usually much bigger than y(0). Explain why most folks are happy to say that cocaine concentration decays exponentially. Not sure how to start

OpenStudy (anonymous):

this is a seperable DE: \[\frac{ dy }{ dt } = \frac{ -ky }{ a+y }\] \[\frac{ (a+y) }{ ky } dy = - dt\] integrate both sides. what school do you go to? O_O

OpenStudy (anonymous):

easier to look at: \[\frac{ (a+y) }{ y } dy = -kdt\]

OpenStudy (anonymous):

let me know if you want the next step(s)

OpenStudy (anonymous):

I'd like to see what happens to the left hand side. Got e^(alogy) . e^y = Ce^-tk How do we get the y on its own?

OpenStudy (anonymous):

I actually don't think you can. my "final answer" looks like: \[e^yy^a = Ce^{-kt}\]

OpenStudy (anonymous):

Ok I'm with you. Good stuff.

OpenStudy (anonymous):

Im studying for a proficiency exam at uiuc and found some old exams online and this was one of them. Im gonna see if i follow but I am a little confused